4) Using the situation in problem 3, lets say M- 1200 kg, vc -20 m/s, m 0.05 kg, and θ = 100 a. What is the initial speed of the apple core? 0.0021 kg/m * v2. What is the e apple core experiences a drag force of Farag- power dissipated by this drag at the moment the apple is released? BONUS: what do you think would happen to such an apple? c. Answers: a) 84.6 km/s b) 1.27 x 1012 w

This is the situation that they are describing.
Answer #1

by conservation of momentum

initial horizontal momentum = final horizontal momentum

1200 * 20 = 1200 * new car velocity * cos(10) + 0.05 * 20

new car velocity = 20.3077 m/s

initial vertical momentum = final vertical momentum

0.05 * apple's velocity = 1200 * 20.3077 * sin(10)

a) apple's velocity = 84633.5 m/s or 84.6 km/s

Fdrag = 0.0021 * v^2

power = force * velocity

power = 0.0021 * v^2 * v

power= 0.0021 * 84633.5^3

b) power = 1.27 * 10^12 W